# Position: Is machine learning good or bad for the natural sciences?

### v1

@@ Abstract
Machine learning (ML) methods are having a huge impact across all of the sciences. However, ML has a strong ontology—in which only the data exist—and a strong epistemology—in which a model is considered good if it performs well on held-out training data. These philosophies are in strong conflict with both standard practices and key philosophies in the natural sciences. Here, we identify some locations for ML in the natural sciences at which the ontology and epistemology are valuable. For example, when an expressive machine learning model is used in a causal inference to represent the effects of confounders, such as foregrounds, backgrounds, or instrument calibration parameters, the model capacity and loose philosophy of ML can make the results more trustworthy. We also show that there are contexts in which the introduction of ML introduces strong, unwanted statistical biases. For one, when ML models are used to emulate physical (or first-principles) simulations, they introduce strong confirmation biases. For another, when expressive regressions are used to label datasets, those labels cannot be used in downstream joint or ensemble analyses without taking on uncontrolled biases. The question in the title is being asked of all of the natural sciences; that is, we are calling on the scientific communities to take a step back and consider the role and value of ML in their fields; the (partial) answers we give here come from the particular perspective of physics.
@@
This definition is broad: In addition to methods like convolutional neural networks (LeCun et al., 1995), multi-layer perceptrons, and transformers (Vaswani et al., 2017), it includes large linear regressions, gaussian processes (Williams & Rasmussen, 2006), support-vector machines (Steinwart & Christmann, 2008), principal components analysis (Greenacre et al., 2022), kernel density estimates (Chen, 2017), and even some kinds of multilayer or hierarchical models (Gelman & Hill, 2006); indeed almost any models in any current ML textbooks (eg, Alpaydin 2020; Acquaviva 2023). Feel free to personally take any more restrictive definition. Our comments will apply to everything in the larger class.
@@
In Bayesian language, this problem is closely related to the point that when data are to be combined, they should be combined at the likelihood level, not the posterior level. Likelihoods are not affected by population-level biases. Likelihoods can produce unbiased point estimates. Likelihood-based estimators are not usually the lowest-risk estimators, but they can be low- or zero-bias estimators. The outputs of regressions, on the other hand, are more like posterior-based estimators, affected by the implicit prior set by the training set. When they are combined, the implicit prior gets amplified. Most ML regressions are not capable of generating or emulating likelihoods. There are interesting approaches in development to replace standard regressions with generative models that can produce likelihoods (Ness et al., 2015; Cranmer et al., 2015; Dai & Seljak, 2022).
@@
Finally, we note that one objection to this confirmation-bias point is that perhaps it exists just as strongly for any simulation-based inference? After all, simulations are expensive, so the full suite of them probably can’t be affordably repeated. There are two responses to this objection. The first is that if the emulators work—that is, if they greatly reduce the cost of the inference—then the emulators have enormously increased the cost difference (or cost factor) between the inference and any re-analysis using physics-based simulations. The second is that emulators are not doing precisely what the simulations did. One way to see this is that they are subject to adversarial attacks (Horowitz & Melchior, 2022). Therefore we know that the emulator is not doing precisely what the physics-based simulation is doing. That is, verification is not just more expensive, but also more important, when an emulator has been introduced.
@@
Given everything we have said here, what are the ways forward for the natural sciences? ML is critical in many projects; how do we make sure it is doing what we want it to do? Our view is that the promising directions include imposing physically motivated mathematical structure into the design or regularization of the ML models. This includes symmetries, conservation laws, and constraints (Bronstein et al., 2021; Villar et al., 2023a), but also ODE/PDE solvers (Karniadakis et al., 2021), dynamic programming modules (Xu et al., 2020), and other established mathematical tools. The research direction that looks towards explainability, interpretability (Molnar, 2020), and trustworthiness (Kearns & Roth, 2019) is also fundamental to the development of ML methods for the physical sciences. These approaches can increase natural scientists’ confidence in ML tools, and also perhaps expand their ontologies and epistemologies.
@@ Acknowledgements:
It is a pleasure to thank Gaby Contardo (SISSA), Jennifer Hill (NYU), Adrian Price-Whelan (Flatiron), Hans-Walter Rix (MPIA), Sam Roweis (deceased), and Bernhard Schölkopf (MPI-IS) for conversations over the years related to these arguments, and Vedant Chandra (Harvard), Contardo, Valentino Foit (NYU), Alex Jiang (CUNY), Charles Margossian (Flatiron), Rix, Kate Storey-Fisher (DIPC), Kaze Wong (Flatiron), and our four anonymous ICML reviewers for valuable comments on this particular manuscript. SV was partially supported by ONR N00014-22-1-2126, NSF CCF 2212457, the NSF–Simons Research Collaboration on the Mathematical and Scientific Foundations of Deep Learning (MoDL) (NSF DMS 2031985), and NSF CAREER 2339682. The Flatiron Institute is a division of the Simons Foundation.

### v2

@@ Abstract
Machine learning (ML) methods are having a huge impact across all of the sciences. However, ML has a strong ontology—in which only the data exist—and a strong epistemology—in which a model is considered good if it performs well on held-out training data. These philosophies are in strong conflict with both standard practices and key philosophies in the natural sciences. Here we identify some locations for ML in the natural sciences at which the ontology and epistemology are valuable. For example, when an expressive machine learning model is used in a causal inference to represent the effects of confounders, such as foregrounds, backgrounds, or instrument calibration parameters, the model capacity and loose philosophy of ML can make the results more trustworthy. We also show that there are contexts in which the introduction of ML introduces strong, unwanted statistical biases. For one, when ML models are used to emulate physical (or first-principles) simulations, they amplify confirmation biases. For another, when expressive regressions are used to label datasets, those labels cannot be used in downstream joint or ensemble analyses without taking on uncontrolled biases. The question in the title is being asked of all of the natural sciences; that is, we are calling on the scientific communities to take a step back and consider the role and value of ML in their fields; the (partial) answers we give here come from the particular perspective of physics.
@@
This definition is broad, perhaps controversially broad: In addition to methods like convolutional neural networks (LeCun et al., 1995), multi-layer perceptrons, and transformers (Vaswani et al., 2017), it includes large linear regressions, Gaussian processes (Williams & Rasmussen, 2006), support-vector machines (Steinwart & Christmann, 2008), principal components analysis (Greenacre et al., 2022), kernel density estimates (Chen, 2017), and even some kinds of multilayer or hierarchical models (Gelman & Hill, 2006); indeed almost any models in any current ML textbooks (eg, Alpaydin 2020; Acquaviva 2023). Feel free to personally take any more restrictive definition. Our comments will apply to everything in the larger class.
@@
In Bayesian language, this problem is closely related to the point that when data are to be combined, they should be combined at the likelihood level, not the posterior level. Likelihoods are not affected by population-level biases. Likelihoods can produce unbiased point estimates. Likelihood-based estimators are not usually the lowest-risk estimators, but they can be low- or zero-bias estimators. The outputs of regressions, on the other hand, are more like posterior-based estimators, affected by the implicit prior set by the training set. When they are combined, the implicit prior gets amplified. Most ML regressions are not capable of generating or emulating likelihoods. There are interesting approaches in development to replace standard regressions with generative models that can produce likelihoods (Ness et al., 2015; Cranmer et al., 2015; Papamakarios et al., 2019; Dai & Seljak, 2022).
@@
Finally, we note that one objection to this confirmation-bias point is that perhaps it exists just as strongly for any simulation-based inference (such as approximate bayesian computation; Csilléry et al. 2010)? After all, simulations are expensive, so the full suite of them probably can’t be affordably repeated. There are two responses to this objection. The first is that if the emulators work—that is, if they greatly reduce the cost of the inference—then the emulators have enormously increased the cost difference (or cost factor) between the inference and any re-analysis using physics-based simulations. The second is that emulators are not doing precisely what the simulations did. One way to see this is that they are subject to adversarial attacks (Horowitz & Melchior, 2022). Therefore we know that the emulator is not doing precisely what the physics-based simulation is doing. That is, verification is not just more expensive, but also more important, when an emulator has been introduced.
@@
Given everything we have said here, what are the ways forward for the natural sciences? ML is critical in many projects; how do we make sure it is doing what we want it to do? Our view is that the promising directions include imposing physically motivated mathematical structure into the design or regularization of the ML models. This includes symmetries, conservation laws, and constraints (Bronstein et al., 2021; Villar et al., 2023a), but also ODE and PDE solvers (Karniadakis et al., 2021), dynamic programming modules (Xu et al., 2020), and other established mathematical tools. The research direction that looks towards explainability, interpretability (Molnar, 2020), and trustworthiness (Kearns & Roth, 2019) is also fundamental to the development of ML methods for the physical sciences. These approaches can increase natural scientists’ confidence in ML tools, and also perhaps expand their ontologies and epistemologies.
@@ Acknowledgements:
It is a pleasure to thank Gaby Contardo (SISSA), Jennifer Hill (NYU), Adrian Price-Whelan (Flatiron), Hans-Walter Rix (MPIA), Sam Roweis (deceased), and Bernhard Schölkopf (MPI-IS) for conversations over the years related to these arguments, and Vedant Chandra (Harvard), Contardo, Valentino Foit (NYU), Alex Jiang (CUNY), Charles Margossian (Flatiron), Rix, Ilya Shpitser (JHU), Kate Storey-Fisher (DIPC), Jeremias Sulam (JHU), Jesse Thaler (MIT), Sebastian Wagner-Carena (Flatiron), Kaze Wong (Flatiron), and our four anonymous ICML reviewers for valuable comments on this particular manuscript. SV was partially supported by ONR N00014-22-1-2126, NSF CCF 2212457, the NSF–Simons Research Collaboration on the Mathematical and Scientific Foundations of Deep Learning (MoDL) (NSF DMS 2031985), and NSF CAREER 2339682. The Flatiron Institute is a division of the Simons Foundation.